# Pullback of arithmetic theta series and its modularity for unitary Shimura curves

### v1

@@ 2.4. Toroidal compactification
$\mathcal{C}_{\Phi}=\underline{\mathrm{Iso}}(\mathcal{L}_{\Phi},\mathcal{O}_{\mathcal{B}_{\Phi}}),\ \mathcal{C}_{\Phi}^{*}=\underline{\mathrm{Hom}}(\mathcal{L}_{\Phi},\mathcal{O}_{\mathcal{B}_{\Phi}}).$
In other words, $\mathcal{C}_{\Phi}^{*}$ is the total space of the line bundle $\mathcal{L}_{\Phi}^{-1}$, and $\mathcal{C}_{\Phi}$ is the complement of the zero section $\mathcal{B}_{\Phi}\hookrightarrow\mathcal{C}_{\Phi}^{*}$.
@@
The group $\Delta_{\Phi}$ acts on $\mathcal{B}_{\Phi}$ by (see [${}^{+}$20, Remark 3.3.2])
@@
The following theorem cites results that we need from [${}^{+}$20, Theorem 3.7.1] which ultimately depends on results from [FC13] and [Lan13].
@@
There is a canonical toroidal compactification $\mathcal{S}\hookrightarrow\mathcal{S}^{*}$ such that $\mathcal{S}^{*}$ is flat over $\mathcal{O}_{\boldsymbol{k}}$ of relative dimension $n-1$. It admits a stratification
@@
[…]
@@
    The Zariski closure of the Cartier divisor $\mathcal{Z}(m)$ in $\mathcal{S}^{*}$ is a Cartier divisor $\mathcal{Z}^{*}(m)$.
  5. (5)
@@
We needs to extend the universal object over $\mathcal{S}$ to $\mathcal{S}^{*}$. The following is due to [How15, Theorem 2.5.2].
@@
The universal abelian scheme $\mathbf{A}$ over $\mathcal{S}$ extends to a semi-abelian group scheme $\mathbf{G}$ over $\mathcal{S}^{*}$ with $\mathcal{O}_{\boldsymbol{k}}$-action such that $\mathbf{G}|_{\mathcal{S}}=\mathbf{A}$. At a geometric point $z=\mathrm{Spec}\,\mathbb{F}\in\mathcal{S}(\Phi)$ where $\Phi=(\mathfrak{m},L)\in\mathrm{Cusp}(L)$ , the semi-abelian group scheme $\mathbf{G}_{z}$ is an extension
@@ 3. Comparison of two line bundles
The goal of this section is to compare two line bundles $\boldsymbol{\omega}$ and $\Omega$ defined respectively in [BHY15] (or [${}^{+}$20]) and [How19], and reinterpret the constant term of the generating series of special divisors (see (3.11) below) in [${}^{+}$20].
We define the line bundle of modular form $\boldsymbol{\omega}$ on $\mathcal{S}$ following [${}^{+}$20, §2.4]. Let $\left(\mathbf{A}_{0},\mathbf{A}\right)$ be the pair of universal abelian schemes over $\mathcal{S}$, let $\mathcal{F}_{\mathbf{A}}\subset\operatorname{Lie}(\mathbf{A})$ be the universal subsheaf of Krämer’s moduli problem. Recall that for any abelian scheme $A\rightarrow S$, we have the following exact sequence of locally free $\mathcal{O}_{\mathcal{S}}$ sheaves
@@
where $H_{1}^{\mathrm{dR}}(A)$ can be defined as the Lie algebra of the universal vector extension of $A$, and $\operatorname{Fil}(A)$ is canonically isomorphic to the $\mathcal{O}_{S}$-dual of $\operatorname{Lie}A^{\vee}$. The principal polarization of $\mathbf{A}$ induces a non-degenerate alternating form $\langle$,$\rangle$ , ⟩ on $H_{1}^{\mathrm{dR}}(\mathbf{A})$, satisfying
@@
Moreover $\mathrm{Fil}(\mathbf{A})$ is totally isotropic with respect to $\langle$,$\rangle$ , ⟩, hence $\langle$,$\rangle$ , ⟩ induces a perfect pairing
@@
be the orthogonal to $\mathcal{F}_{\mathbf{A}}$ under the pairing $\langle$,$\rangle$ , ⟩. It is a rank one $\mathcal{O}_{\mathcal{S}}$-module local direct summand on which $\mathcal{O}_{\boldsymbol{k}}$ acts through the structure morphism $\mathcal{O}_{\boldsymbol{k}}\rightarrow\mathcal{O}_{\mathcal{S}}$. Define the line bundle $\boldsymbol{\omega}$ on $\mathcal{S}$ by
@@
whose kernel $L_{\mathbf{A}}\subset\operatorname{Lie}\mathbf{A}$ is an $\mathcal{O}_{\mathcal{S}}$-module local direct summand of rank one. It is stable under $\mathcal{O}_{\boldsymbol{k}}$, which acts on $\operatorname{Lie}\mathbf{A}/L_{\mathbf{A}}$ and $L_{\mathbf{A}}$ via $\varphi$ and $\bar{\varphi}$ respectively, where $\varphi:\mathcal{O}_{\boldsymbol{k}}\rightarrow\mathcal{O}_{\mathcal{S}}$ is the structure map $\bar{\varphi}$ is its Galois conjugate.
@@
We have the following isomorphism of line bundles on $\mathcal{S}^{*}$.
@@
[…]
(3.5) $\{\epsilon x,\epsilon y\}=\langle\epsilon x,y\rangle.$
[…]
$\{\epsilon x,\epsilon y\}=\langle(\pi\otimes 1)x,y\rangle=-\langle y,(\pi\otimes 1)x\rangle=\langle(\pi\otimes 1)y,x\rangle=\{\epsilon y,\epsilon x\}.$
@@
[…]
@@
$\mathcal{Z}_{p}$ is a regular scheme and a Cartier divisor of $\mathcal{S}_{p}$. Moreover we have the following equation of Cartier divisors.
@@
Let $z$ be a geometric point in $\mathcal{S}_{p}\setminus\mathcal{Z}_{p}$. By the same argument of [Krä03, Theorem 4.5, Step 3], we know that both $\mathrm{Fil}(\mathbf{A}_{z})$ and $\epsilon H_{1}^{\mathrm{dR}}(\mathbf{A}_{z})$ are equal to
$\mathcal{F}_{\mathbf{A}_{z}}^{\bot}\oplus(\epsilon^{-1}\mathcal{F}_{\mathbf{A}_{z}}^{\bot})^{\bot}.$
Here $(\epsilon^{-1}\mathcal{F}_{\mathbf{A}_{z}}^{\bot})^{\bot}$ is the perpendicular complement of $\epsilon^{-1}\mathcal{F}_{\mathbf{A}_{z}}^{\bot}$ in $H_{1}^{dR}(\mathbf{A}_{z})$ with respect to $\langle$,$\rangle$ , ⟩. This implies that $\epsilon=\pi\otimes 1$ acts trivially on $\operatorname{Lie}\mathbf{A}_{z}$, so $z$ lies in $\mathrm{Exc}_{p}$ by definition.
@@
The natural map $L_{\mathbf{A}}\rightarrow\operatorname{Lie}\mathbf{A}/\mathcal{F}_{\mathbf{A}}$ defines a section $s\in\operatorname{Hom}(\Omega^{-1},\boldsymbol{\omega}^{-1})$. We have the following equation on the zero locus of $s$.
@@
Step 1. Let $z$ be a geometric point of $(s)$ with residue field $\kappa$. If the characteristic of $\kappa$ does not divide the discriminant $D$ of $\boldsymbol{k}$, then the structural morphism $\varphi:\mathcal{O}_{\boldsymbol{k}}\rightarrow\mathcal{O}_{\mathcal{S}}$ and its conjugate $\bar{\varphi}$ are not the same. Since $\mathcal{O}_{\boldsymbol{k}}$ acts on $\mathcal{F}_{\mathbf{A}_{z}}$ via $\varphi$, and on $L_{\mathbf{A}_{z}}$ via $\bar{\varphi}$, $L_{\mathbf{A}_{z}}\not\subset\mathcal{F}_{\mathbf{A}_{z}}$, which means the map $s_{z}$ has to be nonzero, a contradiction. Hence $z$ is in $\mathcal{S}_{p}$ for some ramified $p$.
Step 2. Assume $z$ is a geometric point of $\mathcal{S}_{p}$ for a ramified $p$. As $\mathcal{F}_{\mathbf{A}_{z}}^{\bot}\subset\epsilon H_{1}^{\mathrm{dR}}(\mathbf{A}_{z})$, we can assume $\mathcal{F}_{\mathbf{A}_{z}}^{\bot}$ is spanned by $\epsilon x$ for some $x\in H_{1}^{\mathrm{dR}}(\mathbf{A}_{z})$. Notice that $\epsilon^{2}=0$ on $\mathcal{S}_{p}$, hence
$\epsilon^{-1}\mathcal{F}_{\mathbf{A}_{z}}^{\bot}=\mathrm{Span}\{x\}\oplus\epsilon H_{1}^{\mathrm{dR}}(\mathbf{A}_{z}).$
The line $L_{\mathbf{A}_{z}}$ by definition is $\epsilon^{-1}\mathcal{F}_{\mathbf{A}_{z}}^{\bot}/\mathrm{Fil}(\mathbf{A}_{z})\subset\operatorname{Lie}\mathbf{A}_{z}$. Hence
$L_{\mathbf{A}_{z}}\subset\mathcal{F}_{\mathbf{A}_{z}}\Leftrightarrow\langle x,\mathcal{F}_{\mathbf{A}_{z}}^{\bot}\rangle=0\text{ and }\langle\epsilon H_{1}^{\mathrm{dR}}(\mathbf{A}_{z}),\mathcal{F}_{\mathbf{A}_{z}}^{\bot}\rangle=0.$
The last condition is automatic as $\mathcal{F}_{\mathbf{A}_{z}}^{\bot}\subset\epsilon H_{1}^{\mathrm{dR}}(\mathbf{A}_{z})$ by Proposition 3.1 and $\epsilon H_{1}^{\mathrm{dR}}(\mathbf{A}_{z})$ is isotropic. Hence
$L_{\mathbf{A}_{z}}\subset\mathcal{F}_{\mathbf{A}_{z}}\Leftrightarrow\langle x,\mathcal{F}_{\mathbf{A}_{z}}^{\bot}\rangle=0\Leftrightarrow\langle x,\epsilon x\rangle=0.$
[…]
@@
Step 3: We have shown that $(s)$ is supported on the primes ramified in $\boldsymbol{k}$ and its special fiber over such a prime $p$ is $\mathcal{Z}_{p}$. Since $\mathcal{Z}_{p}$ is regular, by [RTZ13, Lemma 10.3], in order to prove the proposition it remains to show that $(s)$ has no $\mathcal{O}_{\boldsymbol{k},p}/(\pi^{2})$-point for any ramified prime $p$. Assume $\tilde{z}$ is such a point. Then on one hand $\pi$ acts on $\mathcal{F}_{\mathbf{A}_{\tilde{z}}}$ by $\varphi(\pi)$. On the other hand $\pi$ acts on $L_{\mathbf{A}_{\tilde{z}}}\subset\mathcal{F}_{\mathbf{A}_{\tilde{z}}}$ by $-\varphi(\pi)$. Since $p\neq 2$, we know that $\pi\neq-\pi$ in $\mathcal{O}_{\boldsymbol{k}}/(\pi^{2})$. This is a contradiction. This finishes the proof of the proposition. ∎

### v2

@@ 2.6. Formal boundary charts
$\mathcal{C}_{\Phi}=\underline{\mathrm{Iso}}(\mathcal{L}_{\Phi},\mathcal{O}_{\mathcal{B}_{\Phi}}),\ \mathcal{C}_{\Phi}^{*}=\underline{\mathrm{Hom}}(\mathcal{L}_{\Phi},\mathcal{O}_{\mathcal{B}_{\Phi}}),$
where $\mathcal{L}_{\Phi}$ is as in (2.12). In other words, $\mathcal{C}_{\Phi}^{*}$ is the total space of the line bundle $\mathcal{L}_{\Phi}^{-1}$, and $\mathcal{C}_{\Phi}$ is the complement of the zero section $\mathcal{B}_{\Phi}\hookrightarrow\mathcal{C}_{\Phi}^{*}$. Relative to $(\mathcal{C}_{\Phi}^{*},\mathcal{B}_{\Phi},\mathcal{C}_{\Phi})$ there is a tautological degeneration data $(B,\kappa,\psi,c,c^{\vee},\tau)$ of type $\Phi$ and signature (n1,1)(n-1,1) where
  • $(B,\kappa,\psi)\in\mathcal{M}(n-2,0)(\mathcal{C}_{\Phi}^{*})$ is as above,
  • $c^{\vee}:\mathfrak{n}\rightarrow B$ is as above and $c=\psi\circ c^{\vee}$,
  • $\tau$ is determined by $\mathcal{C}_{\Phi}^{*}$.
@@
The group $\Delta_{\Phi}$ acts on $\mathcal{B}_{\Phi}$ by (see [BHK+20, Remark 3.3.2])
@@
2.7. Toroidal compactification
For each geometric point $z$ of $\mathcal{B}_{\Phi}$ (viewed as a geometric point of $\mathcal{C}_{\Phi}^{*}$ via the zero section $\mathcal{B}_{\Phi}\hookrightarrow\mathcal{C}_{\Phi}^{*}$), let $R_{z}$ be the étale local ring of $\mathcal{C}_{\Phi}^{*}$ at $z$, and $I_{z}$ be the ideal defined by the divisor $\mathcal{B}_{\Phi}\hookrightarrow\mathcal{C}_{\Phi}^{*}$. Let $\hat{R}_{z}$ be the completion of $R_{z}$ with respect to $I_{z}$, and let $\hat{\eta}_{z}$ be the generic point of $\hat{R}_{z}$. As $\mathcal{C}_{\Phi}^{*}$ is smooth over $\mathcal{O}_{\boldsymbol{k}}$, both $R_{z}$ and $\hat{R}_{z}$ are Noetherian normal domains. By applying Theorem 2.8 to the pullback of the tautological degeneration data relative to $(\mathcal{C}_{\Phi}^{*},\mathcal{B}_{\Phi},\mathcal{C}_{\Phi})$, we get a degenerating abelian scheme $({}^{\heartsuit}G_{z},{}^{\heartsuit}\iota_{z},{}^{\heartsuit}\lambda_{z},{}^{\heartsuit}\mathcal{F}_{z})$ of type $\Phi$ and signature (n1,1)(n-1,1) relative to $(\mathrm{Spec}\,\hat{R_{z}},\mathrm{Spec}\,(\hat{R}_{z}/I_{z}),\hat{\eta}_{z})$. For every étale neighborhood $X^{(z)}\rightarrow\mathcal{C}_{\Phi}^{*}$ of a geometric point $z$, define a closed substack of $X^{(z)}$ by
$Z^{(z)}=\mathcal{B}_{\Phi}\times_{\mathcal{C}_{\Phi}^{*}}X^{(z)},$
and an open substack
$U^{(z)}=\mathcal{C}_{\Phi}\times_{\mathcal{C}_{\Phi}^{*}}X^{(z)}.$
Proposition 2.9.
For every geometric point $z$ of $\mathcal{B}_{\Phi}$ there is an étale neighborhood $X^{(z)}\rightarrow\mathcal{C}_{\Phi}^{*}$ of $z$ and a degenerating abelian scheme $(G^{(z)},\iota^{(z)},\lambda^{(z)},\mathcal{F}^{(z)})$ of type $\Phi$ and signature (n1,1)(n-1,1) relative to $(X^{(z)},Z^{(z)},U^{(z)})$ with the following properties.
  • There exists an automorphism of $\mathcal{O}_{\boldsymbol{k}}$-scheme $\gamma:\mathrm{Spec}\,\hat{R}_{z}\rightarrow\mathrm{Spec}\,\hat{R}_{z}$ inducing the identity on $\mathrm{Spec}\,\hat{R}_{z}/I_{z}$ such that
    $(G^{(z)},\iota^{(z)},\lambda^{(z)},\mathcal{F}^{(z)})|_{\mathrm{Spec}\,\hat{R}_{z}}\cong\gamma^{*}({}^{\heartsuit}G_{z},{}^{\heartsuit}\iota_{z},{}^{\heartsuit}\lambda_{z},{}^{\heartsuit}\mathcal{F}_{z}),$
    where the left hand side is the pullback of $(G^{(z)},\iota^{(z)},\lambda^{(z)},\mathcal{F}^{(z)})$ via the canonical map $\mathrm{Spec}\,(\hat{R}_{z})\rightarrow X^{(z)}$.
  • The tuple $(G^{(z)},\iota^{(z)},\lambda^{(z)},\mathcal{F}^{(z)})|_{U^{(z)}}$ defines an étale morphism
    $U^{(z)}\rightarrow\mathcal{S}.$
  • We have $\mathcal{F}^{(z)}=\mathrm{ker}(\bar{\epsilon}:\operatorname{Lie}G^{(z)}\rightarrow\operatorname{Lie}G^{(z)})$.
Proof.
This is essentially [How15, Proposition 2.5.1] which ultimately depends on results from [Lan13] or [FC13]. ∎
By the quasi-compactness of $\mathcal{B}_{\Phi}$, we may choose finitely many geometric points $z$ so that the union of the images of $X^{(z)}\rightarrow\mathcal{C}_{\Phi}^{*}$ as in Proposition 2.9 covers $\mathcal{B}_{\Phi}$. Let $\Phi$ vary over all cusp labels in $\mathrm{Cusp}([[L]])$, let $\mathcal{X}$ be the disjoint union of the finitely many $X^{(z)}$’s so constructed, and let $\mathcal{U}$ be the disjoint union of the finitely many $U^{(z)}$’s. The obvious map defined by identifying the abelian scheme over $\mathcal{U}$
$\mathcal{S}\sqcup\mathcal{U}\rightarrow\mathcal{S}$
is an étale surjection, and realizes $\mathcal{S}$ as the quotient of $\mathcal{S}\sqcup\mathcal{U}$ by an étale equivalence relation
$\mathcal{R}_{0}\rightarrow(\mathcal{S}\sqcup\mathcal{U})\times_{\mathrm{Spec}\,\mathcal{O}_{\boldsymbol{k}}}(\mathcal{S}\sqcup\mathcal{U}).$
The normalization of $\mathcal{R}_{0}\rightarrow(\mathcal{S}\sqcup\mathcal{X})\times_{\mathrm{Spec}\,\mathcal{O}_{\boldsymbol{k}}}(\mathcal{S}\sqcup\mathcal{X})$ defines a new stack $\mathcal{R}$ sitting in a commutative diagram
(2.18) $\mathcal{R}_{0}$$\mathcal{R}$$(\mathcal{S}\sqcup\mathcal{U})\times_{\mathrm{Spec}\,\mathcal{O}_{\boldsymbol{k}}}(\mathcal{S}\sqcup\mathcal{U})$$(\mathcal{S}\sqcup\mathcal{X})\times_{\mathrm{Spec}\,\mathcal{O}_{\boldsymbol{k}}}(\mathcal{S}\sqcup\mathcal{X}).$$\scriptstyle{r}$
Exactly as in [Lan13, Proposition 6.3.3.13], the morphism $r$ is an étale equivalence relation. Let $\mathcal{S}^{*}$ be the quotient of $\mathcal{S}\sqcup\mathcal{X}$ by $r$. The following theorem cites results that we need from [BHK+20, Theorem 3.7.1].
@@ 2.7. Toroidal compactification
There is a canonical toroidal compactification $\mathcal{S}\hookrightarrow\mathcal{S}^{*}$ such that $\mathcal{S}^{*}$ is flat over $\mathcal{O}_{\boldsymbol{k}}$ of relative dimension n1n-1. It admits a stratification
@@
    $\Delta_{\Phi}\backslash\mathcal{B}_{\Phi}$$\mathcal{S}^{*}(\Phi)$$\Delta_{\Phi}\backslash\mathcal{C}_{\Phi}^{*}$$\mathcal{S}^{*}$$\scriptstyle{\pi_{\Phi}}$
@@
The universal object over $\mathcal{S}$ extends to a semi-abelian scheme over $\mathcal{S}^{*}$. The following is due to [How15, Theorem 2.5.2].
@@
The universal abelian scheme $\mathbf{A}$ over $\mathcal{S}$ extends to a semi-abelian scheme $\mathbf{G}$ over $\mathcal{S}^{*}$ with $\mathcal{O}_{\boldsymbol{k}}$-action such that $\mathbf{G}|_{\mathcal{S}}=\mathbf{A}$. At a geometric point $z=\mathrm{Spec}\,\mathbb{F}\in\mathcal{S}^{*}(\Phi)$ where $\Phi=(\mathfrak{m}\subset M)\in\mathrm{Cusp}([[L]])$ , the semi-abelian scheme $\mathbf{G}_{z}$ is an extension
@@ 3. Comparison of two line bundles
The goal of this section is to compare two line bundles $\boldsymbol{\omega}$ and $\Omega$ defined respectively in [BHY15] (or [BHK+20]) and [How19], and reinterpret the constant term of the generating series of special divisors (see (3.13) below) in [BHK+20].
We define the line bundle of modular form $\boldsymbol{\omega}$ on $\mathcal{S}$ following [BHK+20, §2.4]. Let $\left(\mathbf{A}_{0},\mathbf{A}\right)$ be the pair of universal abelian schemes over $\mathcal{S}$, let $\mathcal{F}_{\mathbf{A}}\subset\operatorname{Lie}(\mathbf{A})$ be the universal subsheaf of Krämer’s moduli problem. Recall that for any abelian scheme $A\rightarrow S$, we have the following exact sequence of locally free $\mathcal{O}_{\mathcal{S}}$ sheaves
@@
where $H_{1}^{\mathrm{dR}}(A)$ can be defined as the Lie algebra of the universal vector extension of $A$, and $\operatorname{Fil}(A)$ is canonically isomorphic to the $\mathcal{O}_{S}$-dual of $\operatorname{Lie}A^{\vee}$. The principal polarization of $\mathbf{A}$ induces a non-degenerate alternating form $\langle,\rangle$ on $H_{1}^{\mathrm{dR}}(\mathbf{A})$, satisfying
@@
Moreover $\mathrm{Fil}(\mathbf{A})$ is totally isotropic with respect to $\langle,\rangle$, hence $\langle,\rangle$ induces a perfect pairing
@@
be the orthogonal to $\mathcal{F}_{\mathbf{A}}$ under the pairing $\langle,\rangle$. It is a rank one $\mathcal{O}_{\mathcal{S}}$-module local direct summand on which $\mathcal{O}_{\boldsymbol{k}}$ acts through the structure morphism $\mathcal{O}_{\boldsymbol{k}}\rightarrow\mathcal{O}_{\mathcal{S}}$. Define the line bundle $\boldsymbol{\omega}$ on $\mathcal{S}$ by
@@
whose kernel $L_{\mathbf{A}}\subset\operatorname{Lie}\mathbf{A}$ is an $\mathcal{O}_{\mathcal{S}}$-module local direct summand of rank one. It is stable under $\mathcal{O}_{\boldsymbol{k}}$, which acts on $\operatorname{Lie}\mathbf{A}/L_{\mathbf{A}}$ and $L_{\mathbf{A}}$ via $\varphi$ and $\bar{\varphi}$ respectively, where $\varphi:\mathcal{O}_{\boldsymbol{k}}\rightarrow\mathcal{O}_{\mathcal{S}}$ is the structure map, and $\bar{\varphi}$ is its Galois conjugate.
@@
We have the following equation in $\mathrm{Pic}_{\mathbb{Q}}(\mathcal{S}^{*})$.
@@
Assume that $p$ is an odd prime of $\mathbb{Q}$ ramified in $\boldsymbol{k}$. Following [Krä03, Theorem 4.5, Step 2], we define a substack $\mathcal{Z}_{p}$ of the special fiber $\mathcal{S}_{p}$ of $\mathcal{S}$ over $p$ as follows. Let $(\mathbf{A}_{0,p},\mathbf{A}_{p})$ be the universal abelian schemes over $\mathcal{S}_{p}$. Set
(3.5) $\varepsilon_{S}=$ $\delta\otimes 1-1\otimes\varphi(\bar{\delta})\in\mathcal{O}_{\boldsymbol{k}}\otimes_{\mathbb{Z}}\mathcal{O}_{S},$
(3.6) $\bar{\varepsilon}_{S}=$ $\bar{\delta}\otimes 1-1\otimes\varphi(\bar{\delta})\in\mathcal{O}_{\boldsymbol{k}}\otimes_{\mathbb{Z}}\mathcal{O}_{S}.$
Notice that $\varepsilon$ induces a morphism from $\varepsilon:H_{1}^{\mathrm{dR}}(\mathbf{A}_{p})\to H_{1}^{\mathrm{dR}}(\mathbf{A}_{p})$ such that the kernel and image of $\varepsilon$ are equal to each other. This is because $H_{1}^{\mathrm{dR}}(\mathbf{A}_{p})$ is locally free over $\mathcal{O}_{\boldsymbol{k}}\otimes\mathcal{S}_{p}$ and the similar property of $\varepsilon$ acting on $\mathcal{O}_{\boldsymbol{k}}\otimes\mathcal{S}_{p}$ holds (on $\mathcal{S}_{p}$ we have $\varepsilon=\delta\otimes 1$). Note that now we have $\varepsilon=-\bar{\varepsilon}$. Define on $\varepsilon H_{1}^{\mathrm{dR}}(\mathbf{A}_{p})$ an $\mathcal{O}_{\mathcal{S}_{p}}$-bilinear form $\{,\}$ by
(3.7) $\{\varepsilon x,\varepsilon y\}=\langle\varepsilon x,y\rangle.$
As $\varepsilon H_{1}^{\mathrm{dR}}(\mathbf{A}_{p})$ is isotropic with respect to $\langle,\rangle$, the form $\{,\}$ is well-defined. Moreover $\{,\}$ is symmetric:
$\{\varepsilon x,\varepsilon y\}=\langle(\pi\otimes 1)x,y\rangle=-\langle y,(\pi\otimes 1)x\rangle=\langle(\pi\otimes 1)y,x\rangle=\{\varepsilon y,\varepsilon x\}.$
@@
Let $\mathcal{Z}_{p}$ be the substack of $\mathcal{S}_{p}$ such that for any $\mathcal{O}_{\boldsymbol{k}}/(\mathfrak{p})$-scheme $S$, $\mathcal{Z}_{p}(S)$ is the groupoid of isomorphism classes $(A_{0},\iota_{0},\lambda_{0},A,\iota,\lambda,\mathcal{F})\in\mathcal{S}_{p}(S)$ such that $\mathcal{F}_{A}^{\bot}$ is isotropic with respect to $\{,\}$. Here $\mathfrak{p}$ is the ideal of $\mathcal{O}_{\boldsymbol{k}}$ such that $\mathfrak{p}^{2}=p\mathcal{O}_{\boldsymbol{k}}$.
@@
$\mathcal{Z}_{p}$ is a regular scheme and a Cartier divisor of $\mathcal{S}$. Moreover we have the following equation of Cartier divisors
@@
Let $z$ be a geometric point in $\mathcal{S}_{p}\setminus\mathcal{Z}_{p}$. By the same argument of [Krä03, Theorem 4.5, Step 3], we know that both $\mathrm{Fil}(\mathbf{A}_{z})$ and $\varepsilon H_{1}^{\mathrm{dR}}(\mathbf{A}_{z})$ are equal to
$\mathcal{F}_{\mathbf{A}_{z}}^{\bot}\oplus(\varepsilon^{-1}\mathcal{F}_{\mathbf{A}_{z}}^{\bot})^{\bot}.$
Here $(\varepsilon^{-1}\mathcal{F}_{\mathbf{A}_{z}}^{\bot})^{\bot}$ is the perpendicular complement of $\varepsilon^{-1}\mathcal{F}_{\mathbf{A}_{z}}^{\bot}$ in $H_{1}^{dR}(\mathbf{A}_{z})$ with respect to $\langle,\rangle$. This implies that $\varepsilon=\pi\otimes 1$ acts trivially on $\operatorname{Lie}\mathbf{A}_{z}$, so $z$ lies in $\mathrm{Exc}_{p}$ by definition.
@@
The natural map $L_{\mathbf{A}}\rightarrow\operatorname{Lie}\mathbf{A}/\mathcal{F}_{\mathbf{A}}$ defines a section $s\in\operatorname{Hom}(\Omega^{-1},\boldsymbol{\omega}^{-1})$. We have the following equation for the zero locus of $s$.
@@
Step 1. Let $z$ be a geometric point of $(s)$ with residue field $\kappa$. If the characteristic of $\kappa$ does not divide the discriminant $D$ of $\boldsymbol{k}$, then the structural morphism $\varphi:\mathcal{O}_{\boldsymbol{k}}\rightarrow\kappa$ and its conjugate $\bar{\varphi}$ are not the same. Since $\mathcal{O}_{\boldsymbol{k}}$ acts on $\mathcal{F}_{\mathbf{A}_{z}}$ via $\varphi$, and on $L_{\mathbf{A}_{z}}$ via $\bar{\varphi}$, $L_{\mathbf{A}_{z}}\not\subset\mathcal{F}_{\mathbf{A}_{z}}$, which means the map $s_{z}$ has to be nonzero, a contradiction. Hence $z$ is in $\mathcal{S}_{p}$ for some ramified $p$.
Step 2. Assume $z$ is a geometric point of $\mathcal{S}_{p}$ for a ramified $p$. As $\mathcal{F}_{\mathbf{A}_{z}}^{\bot}\subset\varepsilon H_{1}^{\mathrm{dR}}(\mathbf{A}_{z})$, we can assume $\mathcal{F}_{\mathbf{A}_{z}}^{\bot}$ is spanned by $\varepsilon x$ for some $x\in H_{1}^{\mathrm{dR}}(\mathbf{A}_{z})$. Since the kernel and image of $\varepsilon$ acting on $H_{1}^{\mathrm{dR}}(\mathbf{A}_{z})$ are equal to each other, we have
$\varepsilon^{-1}\mathcal{F}_{\mathbf{A}_{z}}^{\bot}=\mathrm{Span}\{x\}\oplus\varepsilon H_{1}^{\mathrm{dR}}(\mathbf{A}_{z}).$
The line $L_{\mathbf{A}_{z}}$ by definition is $\varepsilon^{-1}\mathcal{F}_{\mathbf{A}_{z}}^{\bot}/\mathrm{Fil}(\mathbf{A}_{z})\subset\operatorname{Lie}\mathbf{A}_{z}$. Hence
$L_{\mathbf{A}_{z}}\subset\mathcal{F}_{\mathbf{A}_{z}}\Leftrightarrow\langle x,\mathcal{F}_{\mathbf{A}_{z}}^{\bot}\rangle=0\text{ and }\langle\varepsilon H_{1}^{\mathrm{dR}}(\mathbf{A}_{z}),\mathcal{F}_{\mathbf{A}_{z}}^{\bot}\rangle=0.$
The last condition is automatic as $\mathcal{F}_{\mathbf{A}_{z}}^{\bot}\subset\varepsilon H_{1}^{\mathrm{dR}}(\mathbf{A}_{z})$ by Proposition 3.1 and $\varepsilon H_{1}^{\mathrm{dR}}(\mathbf{A}_{z})$ is isotropic. Hence
$L_{\mathbf{A}_{z}}\subset\mathcal{F}_{\mathbf{A}_{z}}\Leftrightarrow\langle x,\mathcal{F}_{\mathbf{A}_{z}}^{\bot}\rangle=0\Leftrightarrow\langle x,\varepsilon x\rangle=0.$
By the definition of $\{,\}$, the last condition is true if and only if $\{x,x\}=0$, i.e., $\mathcal{F}_{\mathbf{A}_{z}}^{\bot}$ is isotropic with respect to $\{,\}$. Hence
@@
Step 3: We have shown that $(s)$ is supported on the primes ramified in $\boldsymbol{k}$ and its special fiber over such a prime $p$ is $\mathcal{Z}_{p}$. Since $\mathcal{Z}_{p}$ is regular, by [RTZ13, Lemma 10.3], in order to prove the proposition it remains to show that $(s)$ has no $\mathcal{O}_{\boldsymbol{k},p}/(\pi^{2})$-point for any ramified prime $p$. Assume $\tilde{z}$ is such a point. Then on one hand $\pi$ acts on $\mathcal{F}_{\mathbf{A}_{\tilde{z}}}$ by $\varphi(\pi)$. On the other hand $\pi$ acts on $L_{\mathbf{A}_{\tilde{z}}}\subset\mathcal{F}_{\mathbf{A}_{\tilde{z}}}$ by $\varphi(\bar{\pi})$. Since $p\neq 2$, we know that $\pi\neq\bar{\pi}$ in $\mathcal{O}_{\boldsymbol{k}}/(\pi^{2})$. This is a contradiction. This finishes the proof of the proposition. ∎